OpenStudy (gelow):

Ben bought a digital camera for $90. The tax rate is 6%. Which proportion can be used to find the amount of tax Ben will pay? A. Working with Percent B. Working with Percent C. Working with Percent D. Working with Percent

OpenStudy (gelow):

hold on

OpenStudy (gelow):

A x/90 = 6/100 b x/6 = 100/90 c 6/x =90/100 d 90/x =6/100

OpenStudy (gelow):

but the answers are A x/90 = 6/100 b x/6 = 100/90 c 6/x =90/100 d 90/x =6/100

OpenStudy (gelow):

@Iloveyou;)boy